# DebugAgent: Efficient and Interpretable Error Slice Discovery for Comprehensive Model Debugging

### v1

@@ Abstract
Despite the significant success of deep learning models in computer vision, they often exhibit systematic failures on specific data subsets, known as error slices. Identifying and mitigating these error slices is crucial to enhancing model robustness and reliability in real-world scenarios. In this paper, we introduce DebugAgent, an automated framework for error slice discovery and model repair. DebugAgent first generates task-specific visual attributes to highlight instances prone to errors through an interpretable and structured process. It then employs an efficient slice enumeration algorithm to systematically identify error slices, overcoming the combinatorial challenges that arise during slice exploration. Additionally, DebugAgent extends its capabilities by predicting error slices beyond the validation set, addressing a key limitation of prior approaches. Extensive experiments across multiple domains — including image classification, pose estimation, and object detection — show that DebugAgent not only improves the coherence and precision of identified error slices but also significantly enhances the model repair capabilities. https://github.com/cure-lab/DebugAgent.
@@ 1 Introduction
In this paper, we introduce DebugAgent, a fully automated, closed-loop debug framework. As depicted in Figure 1, DebugAgent encompasses attribute and tag generation, error slice discovery, and model repair. To generate comprehensive visual attributes, we implement a structured generation process informed by model failure analysis and engineering insights. To mitigate the combinatorial explosion issue, we develop an efficient slice enumeration algorithm based on the unique characteristics of data slices, along with slice selection and image querying techniques to facilitate model repair. Additionally, we also employ feature-based tag substitutions and instruction-based methods to address unseen errors beyond the validation set.
Our experiments across image classification, pose estimation, and object detection tasks, spanning multiple datasets and models, demonstrate the superior performance of DebugAgent. Specifically, DebugAgent consistently produces attributes of significantly higher quality than existing methods, while its slice enumeration algorithm achieves an impressive 510x speedup over naive approaches. Furthermore, DebugAgent shows strong generalizability in identifying error slices on widely-used models. For instance, we identified approximately 500 distinct error slices for CLIP in image classification tasks. In addition, the classification models involved in the experiments exhibit performance declines of up to 64.6% on predicted unseen error slices. Finally, experimental results validate that DebugAgent outperforms prior methods in its model repair capabilities.
Figure 1: The main workflow of DebugAgent for closed-loop model debugging and repair.
@@ 3 Method: automatic attribute and tag generation
We present the workflow of DebugAgent in Figure 1. Attribute and tag generation serves as the foundation of DebugAgent, as it is closely linked to the coherence and coverage of error slices. Our attribute and tag generation consists of the following steps: attribute generation, tag determination, and dataset-wide tag assignment. It addresses several key challenges identified in existing approaches.
@@ 3.2 Structured and comprehensive generation
To address these challenges, DebugAgent leverages the strengths of multi-modal models (in this paper, we use GPT (OpenAI, 2023)) for attribute and tag generation, combined with a structured process to ensure the accuracy, consistency, and coverage of the results.
@@ 3.2.2 Tag determination and assignment
DebugAgent begins by generating an initial list of potential and unbiased tags for each attribute, establishing the semantic scope and granularity. To ensure that the generated tags are concrete and unambiguous, we employ clear conventions: for binary attributes, such as “whether an object is occluded”, we adopt a straightforward “yes/no” tag format, while for open-ended attributes, DebugAgent generates a descriptive set of tags. Next, DebugAgent collects new tags by reviewing a subset of validation data, incorporating additional tags as needed to ensure comprehensive coverage of variations. Once the attributes and tags are finalized, DebugAgent assigns these tags to all images within the dataset, ensuring that each image receives a tag corresponding to each attribute. This multi-stage process promotes consistency in tags throughout the dataset, thereby enhancing the robustness of subsequent analyses. We discuss DebugAgent’s scalability and potential issues regrading attribute and tag generation in Appendix A.8.
@@ 4.1.4 Algorithm design
Breadth-First Tree-structured Enumeration. To address the above challenges, we propose a breadth-first tree-structured enumeration process. We establish a tree based on the generated attributes and tags, in which the slices stored in child nodes are formed by adding one attribute-tag pair to the slices stored in their parent nodes. The parent-child relationships of slices can then be represented by the parent-child relationships of nodes, and the numbers of attributes in the slices grow with the tree’s depth. Note that a child node may correspond to multiple parent nodes in the tree. Crucially, since the data count monotonically decreases with deeper layers, data enumeration of a slice can be upper-bounded by its parent slices, significantly reducing the search space. We employ breadth-first search (Bellman, 1958) (BFS) to ensure that parent slices are enumerated before child slices.
@@
Integration with Model Repair. We incorporate image querying techniques for model repair. After post-processing, DebugAgent first ranks the slices based on their average model performance. Given a data pool, DebugAgent then assigns tags to the data and prioritizes those corresponding to error slices with the lowest average performance. This ensures that the most critical cases are addressed first, effectively targeting the model’s weakest points for repair.
@@ 4.2 Predicting error slices beyond the validation set
Instruction-Based Method. We utilize few-shot learning with GPT to predict potential error slices based on the provided attributes and tags, following task-specific instructions. These instructions are similar to those used for generating error-related attributes. For instance, in image classification tasks, we instruct the model to generate slices that blur boundaries between closely related categories. In object detection and pose estimation tasks, we focus on generating slices prone to localization errors or occlusions. Unlike Tag Substitution which is built upon the model’s identified error slices, the instructions-based method leverages GPT’s extensive knowledge base without any prior information about the model’s errors.
@@ 5.2 Effectiveness of slice enumeration
As shown in Figure 5, our method achieves approximately 115x and 510x speedups over naive enumeration, and 12x and 7x speedups compared to the baseline tree-structured algorithm, for the enumeration of slices with 3 and 4 attributes respectively. These significant improvements in efficiency allow for rapid slice enumeration across multiple attributes, substantially reducing the computational time required for model analysis in real-world scenarios. Additionally, we conduct an ablation study focusing on the enumeration of slices with 3 attributes to assess the effect of the data volume and total number of attributes on runtime. The results in Figure 5 demonstrate that runtime increases linearly with the data volume, while remaining feasible even in cases where tasks involve up to 72 attributes.
@@ 5.3 Identified error slices
We conduct experiments across three tasks — image classification, pose estimation, and object detection — to evaluate our method’s ability to identify error slices. For image classification, we select five bear species from ImageNet (Deng et al., 2009) and debug three models: ResNet18 (He et al., 2016), CLIP (Radford et al., 2021), and BLIP (Li et al., 2022a). For pose estimation, we use an industrial private dataset and debug the tiny, small, medium, and large variants of RTMPose (Jiang et al., 2023). For object detection, we use the “Car” and the “Pedestrian” instances of the KITTI (Geiger et al., 2012) dataset and debug four models: YOLOv8 (Varghese & Sambath, 2024), CO-DINO (Zong et al., 2023), ViTDet-L (Li et al., 2022b), and RTMDet-X (Lyu et al., 2022). We apply DebugAgent to automatically generate attributes and tags and perform slice enumeration, followed by error slice analysis based on the results. In the main paper, we consider slices with three attributes. Implementation details of all models and datasets are shown in Appendix A.3.
Figure 6: Identified slices of the image classification task by DebugAgent.
Figure 7: Identified slices of the object detection task by DebugAgent.
@@ 5.5 Model repair
In this section, we evaluate the model repair capabilities across the three previously discussed tasks, focusing on querying new data based on the identified error slices for model improvement. For each task, we construct a query set from which additional data is selected, and the model is evaluated on a hold-out test set, both distinct from the validation set used for identifying error slices. We compare DebugAgent with HiBug (Chen et al., 2024), which is also an automated method, as well as random data selection. For a fair comparison, we implement the same data selection strategy for both DebugAgent and HiBug, prioritizing data that corresponds to slices with the lowest average model performance. The model we repair for the three tasks are respectively ResNet18, RTMPose-Tiny, and RTMDet-X. Implementation details can be found in Appendix A.5.
We summarize the model’s improvements in terms of accuracy, keypoint average precision (AP), and object mean average precision (mAP) for the three tasks in Table 2. DebugAgent consistently outperforms other methods; notably, it enhances model performance when random data selection yields only marginal improvements. This underscores its effectiveness in model repair.
Table 2: Comparisons of model improvements with data slices determined by DebugAgent, HiBug, and random selection. The values are averaged over five runs.
@@ 6 Discussion
Our experiments demonstrate that DebugAgent significantly advances error slice discovery. It improves both the coherence and coverage of identified data slices, which leads to a more interpretable and insightful error analysis process. Meanwhile, the efficient slice enumeration algorithm allows for the rapid discovery of slices across multiple attributes, enabling a more granular analysis of model errors.
However, there are some limitations associated with DebugAgent. The attribute and tag generation process relies on GPT, which can occasionally produce errors. A primary concern might be the impact of incorrect tag assignments. Though such errors might slightly affect the coherence of error slices, they are unlikely to influence overall identification, as a few misclassified data points do not alter the average performance of a slice. We further discuss several potential issues and the scalability of DebugAgent in Appendix A.8.
Moreover, we observe that existing slice discovery methods often follow diverse workflows, complicating direct and equitable comparisons. Future work can focus on developing a standardized and widely applicable benchmark to facilitate fair evaluations and drive progress in this area.
@@ 7 Conclusion
In this paper, we introduce DebugAgent, a comprehensive framework for efficient and interpretable error slice discovery aimed at enhancing model debugging and repair. By leveraging task-specific attribute generation, an efficient slice enumeration algorithm, and prediction of unseen error slices, DebugAgent significantly improves the coherence and coverage of identified slices across diverse tasks, including image classification, pose estimation, and object detection. Our extensive experiments demonstrate that DebugAgent not only outperforms existing methods in terms of model repair and error slice discovery but also provides deeper insights into model failures. We believe DebugAgent’s capabilities can drive broader adoption of slice-based debugging techniques in both academic and industrial settings.
@@ Appendix A Appendix
A.1 Scalability and robustness of DebugAgent
We address potential concerns regarding the scalability and robustness of DebugAgent, particularly focusing on the attribute and tag generation process. We hope that this discussion provides additional clarity and insight into the scalability and robustness of DebugAgent. We will also add alternative versions of DebugAgent with reliance on other multi-modal models, such as LLaVA (Liu et al., 2024) and QWen-VL (Bai et al., 2023), in the future.
@@ A.1.3 Handling Unnecessary Attributes:
It is difficult to assess the necessity of specific attributes prior to the error slice discovery phase. DebugAgent generates a substantially larger set of attributes than existing methods, some of which may not directly align with human recognition. However, only after analysis can the utility of these attributes be determined. For instance, in our classification analysis using BLIP, an attribute such as “is man-made object presented” may initially appear irrelevant. Yet, error slice discovery reveals that BLIP tends to misclassify black bears as sloth bears when man-made objects, such as the iron fence of the zoo, is present in the image. Furthermore, we have developed a post-processing algorithm that effectively removes unnecessary attributes during the error slice discovery stage, ensuring that only the relevant attributes are retained for slice analysis.
@@ A.1.4 Generalization to Different Tasks and Datasets:
Extending DebugAgent to other tasks primarily involves adjusting the attribute and tag generation process, particularly the task-specific prompts used for generating relevant attributes. We provide prompt templates for easy task extension, allowing users to extend our method by simply modifying task descriptions. For different datasets within the same task, the primary variation lies in the change of the main object. Through additional experiments on the full ImageNet and COCO detection datasets (not included in the paper), we observe that DebugAgent generalizes well across a wide range of objects, demonstrating its scalability and flexibility for various tasks and datasets.
A.1.5 Time cost of DebugAgent.
The most time-consuming part of DebugAgent is the process of tag assignment for all the images in the dataset, which typically accounts for over 95% of the total time, with larger datasets increasing this proportion. It operates with $O(N)$, where $N$ is the number of images in a dataset, because GPT-4V generate all attributes for one image in one step. In our experiments, a single query takes around 6 seconds. We implement parallel processing to accelerate these processes, since the tag assignments for different data are independent. For a dataset of 10,000 images, if we use 50 threads to accelerate the labeling process, it typically takes around 30 minutes to run DebugAgent. It is also worth noting that attribute and tag generation and slice enumeration only need to be performed once for a dataset. When iteratively improving the performance of a model, only the first round requires labeling and slice enumeration. Subsequent iterations only require the post-processing steps described in Section 4.1.4, which takes only a few seconds.
@@ A.1.7 Are error slices distinct bugs?
A.2 Generated attributes and tags by DebugAgent and the baseline methods
@@
Table 5: The full list of attributes and tags generated by DebugAgent for the pose estimation task. For tag sets containing more than five tags, the remaining tags are omitted with ellipses
@@
We present a full list of attributes and tags generated by DebugAgent and HiBug for pose estimation in Table 4 and Table 5. The attributes generated by HiBug primarily focus on the main object, often neglecting task-specific requirements and potential errors. Furthermore, the tags frequently lack semantic consistency; for example, the tags associated with the attribute “age” include overlapping terms such as “young”, “teen”, and “12”. Similarly, the tags for “skin tone” combine both race and color categories, resulting in semantic ambiguity. This inconsistency can adversely affect the data slicing process, leading to images with similar semantics being assigned to different slices (e.g., slice “teen” versus slice “young”). In contrast, the attributes and tags generated by DebugAgent are specifically tailored to tasks and errors, ensuring semantic consistency. This makes DebugAgent more effective for model debugging and refinement.

### v3

# HiBug2: Efficient and Interpretable Error Slice Discovery for Comprehensive Model Debugging

@@ Abstract
Despite the significant success of deep learning models in computer vision, they often exhibit systematic failures on specific data subsets, known as error slices. Identifying and mitigating these error slices is crucial to enhancing model robustness and reliability in real-world scenarios. In this paper, we introduce HiBug2, an automated framework for error slice discovery and model repair. HiBug2 first generates task-specific visual attributes to highlight instances prone to errors through an interpretable and structured process. It then employs an efficient slice enumeration algorithm to systematically identify error slices, overcoming the combinatorial challenges that arise during slice exploration. Additionally, HiBug2 extends its capabilities by predicting error slices beyond the validation set, addressing a key limitation of prior approaches. Extensive experiments across multiple domains — including image classification, pose estimation, and object detection — show that HiBug2 not only improves the coherence and precision of identified error slices but also significantly enhances the model repair capabilities. Our code is available at https://github.com/cure-lab/HiBug2.
@@ 1 Introduction
In this paper, we introduce HiBug2, a fully automated, closed-loop debug framework. As depicted in Figure 1, HiBug2 encompasses attribute and tag generation, error slice discovery, and model repair. To generate comprehensive visual attributes, we implement a structured generation process informed by model failure analysis and engineering insights. To mitigate the combinatorial explosion issue, we develop an efficient slice enumeration algorithm based on the unique characteristics of data slices, along with slice selection and image querying techniques to facilitate model repair. Additionally, we also employ feature-based tag substitutions and instruction-based methods to address unseen errors beyond the validation set.
Our experiments across image classification, pose estimation, and object detection tasks, spanning multiple datasets and models, demonstrate the superior performance of HiBug2. Specifically, HiBug2 consistently produces attributes of significantly higher quality than existing methods, while its slice enumeration algorithm achieves an impressive 510x speedup over naive approaches. Furthermore, HiBug2 shows strong generalizability in identifying error slices on widely-used models. For instance, we identified approximately 500 distinct error slices for CLIP in image classification tasks. In addition, the classification models involved in the experiments exhibit performance declines of up to 64.6% on predicted unseen error slices. Finally, experimental results validate that HiBug2 outperforms prior methods in its model repair capabilities.
@@ 2.1 Error slice discovery
Figure 1: The main workflow of HiBug2 for closed-loop model debugging and repair.
@@ 3 Method: automatic attribute and tag generation
We present the workflow of HiBug2 in Figure 1. Attribute and tag generation serves as the foundation of HiBug2, as it is closely linked to the coherence and coverage of error slices. Our attribute and tag generation consists of the following steps: attribute generation, tag determination, and dataset-wide tag assignment. It addresses several key challenges identified in existing approaches.
@@ 3.2 Structured and comprehensive generation
To address these challenges, HiBug2 leverages the strengths of multi-modal models (in this paper, we use GPT (OpenAI, 2023)) for attribute and tag generation, combined with a structured process to ensure the accuracy, consistency, and coverage of the results.
@@ 3.2.2 Tag determination and assignment
HiBug2 begins by generating an initial list of potential and unbiased tags for each attribute, establishing the semantic scope and granularity. To ensure that the generated tags are concrete and unambiguous, we employ clear conventions: for binary attributes, such as “whether an object is occluded”, we adopt a straightforward “yes/no” tag format, while for open-ended attributes, HiBug2 generates a descriptive set of tags. Next, HiBug2 collects new tags by reviewing a subset of validation data, incorporating additional tags as needed to ensure comprehensive coverage of variations. Once the attributes and tags are finalized, HiBug2 assigns these tags to all images within the dataset, ensuring that each image receives a tag corresponding to each attribute. This multi-stage process promotes consistency in tags throughout the dataset, thereby enhancing the robustness of subsequent analyses. We discuss HiBug2’s scalability and potential issues regrading attribute and tag generation in Appendix A.8.
@@ 4.1.4 Algorithm design
Breadth-First Tree-structured Enumeration. To address the above challenges, we propose a breadth-first tree-structured enumeration process. We establish a tree based on the generated attributes and tags, in which slices stored in child nodes are formed by adding one attribute-tag pair to slices stored in their parent nodes. The parent-child relationships of slices can then be represented by the parent-child relationships of nodes, and the numbers of attributes in the slices grow with the tree’s depth. Note that a child node may correspond to multiple parent nodes in the tree. Crucially, since the data count monotonically decreases with deeper layers, data enumeration of a slice can be upper-bounded by its parent slices, significantly reducing the search space. We employ breadth-first search (Bellman, 1958) (BFS) to ensure that parent slices are enumerated before child slices.
@@
Integration with Model Repair. We incorporate image querying techniques for model repair. After post-processing, HiBug2 first ranks the slices based on their average model performance. Given a data pool, HiBug2 then assigns tags to the data and prioritizes those corresponding to error slices with the lowest average performance. This ensures that the most critical cases are addressed first, effectively targeting the model’s weakest points for repair.
@@ 4.2 Predicting error slices beyond the validation set
Instruction-Based Method. We utilize few-shot learning with GPT to predict potential error slices based on the provided attributes and tags, following task-specific instructions. These instructions are similar to those used for generating error-related attributes. For instance, in image classification, we instruct the model to generate slices that blur boundaries between closely related categories. In object detection and pose estimation, we focus on generating slices prone to localization errors or occlusions. Unlike Tag Substitution which is built upon identified error slices, the instructions-based method leverages GPT’s extensive knowledge base without any prior information about the model.
@@ 5.2 Effectiveness of slice enumeration
As shown in Figure 5, our method achieves approximately 115x and 510x speedups over naive enumeration, and 12x and 7x speedups compared to the baseline tree-structured algorithm, for the enumeration of slices with 3 and 4 attributes respectively. These significant improvements in efficiency allow for rapid slice enumeration across multiple attributes, substantially reducing the computational time required for model analysis in real-world scenarios. Additionally, we conduct ablation studies focusing on the enumeration of slices with 3 attributes to assess the effect of data volume and total number of attributes in runtime. The results in Figure 5 demonstrate that runtime increases linearly with the data volume, while remaining feasible even in cases where tasks involve up to 72 attributes.
@@ 5.3 Identified error slices
We conduct experiments across three tasks — image classification, pose estimation, and object detection — to evaluate our method’s ability to identify error slices. For image classification, we select five bear species from ImageNet (Deng et al., 2009) and debug three models: ResNet18 (He et al., 2016), CLIP (Radford et al., 2021), and BLIP (Li et al., 2022a). For pose estimation, we use an industrial private dataset and debug the tiny, small, medium, and large variants of RTMPose (Jiang et al., 2023). For object detection, we use the “Car” and the “Pedestrian” instances of the KITTI (Geiger et al., 2012) dataset and debug four models: YOLOv8 (Varghese & Sambath, 2024), CO-DINO (Zong et al., 2023), ViTDet-L (Li et al., 2022b), and RTMDet-X (Lyu et al., 2022). We apply HiBug2 to automatically generate attributes and tags and perform slice enumeration, followed by error slice analysis based on the results. In the main paper, we consider slices with three attributes. Implementation details of all models and datasets are shown in Appendix A.3.
Figure 6: Identified slices of the image classification task by HiBug2.
Figure 7: Identified slices of the object detection task by HiBug2.
@@ 5.5 Model repair
In this section, we evaluate the model repair capabilities across the three previously discussed tasks, focusing on querying new data based on the identified error slices for model improvement. For each task, we construct a query set from which additional data is selected, and the model is evaluated on a hold-out test set, both distinct from the validation set used for identifying error slices. We compare HiBug2 with HiBug (Chen et al., 2024), which is also an automated method, as well as random data selection. For a fair comparison, we implement the same data selection strategy for both HiBug2 and HiBug, prioritizing data that corresponds to slices with the lowest average model performance. The model we repair for the three tasks are respectively ResNet18, RTMPose-Tiny, and RTMDet-X. Implementation details can be found in Appendix A.5.
We summarize the model’s improvements in terms of accuracy, keypoint average precision (AP), and object mean average precision (mAP) for the three tasks in Table 2. HiBug2 consistently outperforms other methods; notably, it enhances model performance when random data selection yields only marginal improvements. This underscores its effectiveness in model repair.
Table 2: Comparisons of model improvements with data slices determined by HiBug2, HiBug, and random selection. The values are averaged over five runs.
@@ 6 Discussion
Our experiments demonstrate that HiBug2 significantly advances error slice discovery. It improves both coherence and coverage of identified data slices, which leads to a more interpretable and insightful error analysis process. Meanwhile, the efficient slice enumeration algorithm allows for rapid discovery of slices across multiple attributes, enabling a more granular analysis of model errors.
However, there are some limitations associated with HiBug2. Attribute and tag generation leverages GPT that may occasionally produce errors. A primary concern might be the impact of incorrect tag assignments. Though such errors may slightly affect the coherence of error slices, they are unlikely to influence overall identification, as a few misclassified data points do not alter the average performance of a slice. We further discuss several potential issues and the scalability of HiBug2 in Appendix A.8. Moreover, we observe that existing slice discovery methods follow diverse workflows, complicating direct and equitable comparisons. Future work can focus on developing a standardized and widely applicable benchmark to facilitate fair evaluations and drive progress in this area.
@@ 7 Conclusion
In this paper, we introduce HiBug2, a comprehensive framework for efficient and interpretable error slice discovery aimed at enhancing model debugging and repair. By leveraging task-specific attribute generation, efficient slice enumeration, and prediction of unseen error slices, HiBug2 significantly improves coherence and coverage of identified slices across diverse tasks, including image classification, pose estimation, and object detection. Our extensive experiments demonstrate that HiBug2 not only outperforms existing methods in terms of error slice discovery and model repair but also provides deeper insights into model failures. We believe HiBug2’s capabilities can drive broader adoption of slice-based debugging techniques in both academic and industrial settings.
8 Acknowledgments
This work was supported in part by the CUHK Strategic Seed Funding for Collaborative Research Scheme under Grant No. 3136023.
@@ Appendix A Appendix
A.1 Scalability and robustness of HiBug2
We address potential concerns regarding the scalability and robustness of HiBug2, particularly focusing on the attribute and tag generation process. We hope that this discussion provides additional clarity and insight into the scalability and robustness of HiBug2. We will also add alternative versions of HiBug2 with reliance on other multi-modal models, such as LLaVA (Liu et al., 2024) and QWen-VL (Bai et al., 2023), in the future.
@@ A.1.3 Handling Unnecessary Attributes:
It is difficult to assess the necessity of specific attributes prior to the error slice discovery phase. HiBug2 generates a substantially larger set of attributes than existing methods, some of which may not directly align with human recognition. However, only after analysis can the utility of these attributes be determined. For instance, in our classification analysis using BLIP, an attribute such as “is man-made object presented” may initially appear irrelevant. Yet, error slice discovery reveals that BLIP tends to misclassify black bears as sloth bears when man-made objects, such as the iron fence of the zoo, is present in the image. Furthermore, we have developed a post-processing algorithm that effectively removes unnecessary attributes during the error slice discovery stage, ensuring that only the relevant attributes are retained for slice analysis.
@@ A.1.4 Generalization to Different Tasks and Datasets:
Extending HiBug2 to other tasks primarily involves adjusting the attribute and tag generation process, particularly the task-specific prompts used for generating relevant attributes. We provide prompt templates for easy task extension, allowing users to extend our method by simply modifying task descriptions. For different datasets within the same task, the primary variation lies in the change of the main object. Through additional experiments on the full ImageNet and COCO detection datasets (not included in the paper), we observe that HiBug2 generalizes well across a wide range of objects, demonstrating its scalability and flexibility for various tasks and datasets.
A.1.5 Time cost of HiBug2.
The most time-consuming part of HiBug2 is the process of tag assignment for all the images in the dataset, which typically accounts for over 95% of the total time, with larger datasets increasing this proportion. It operates with $O(N)$, where $N$ is the number of images in a dataset, because GPT-4V generate all attributes for one image in one step. In our experiments, a single query takes around 6 seconds. We implement parallel processing to accelerate these processes, since the tag assignments for different data are independent. For a dataset of 10,000 images, if we use 50 threads to accelerate the labeling process, it typically takes around 30 minutes to run HiBug2. It is also worth noting that attribute and tag generation and slice enumeration only need to be performed once for a dataset. When iteratively improving the performance of a model, only the first round requires labeling and slice enumeration. Subsequent iterations only require the post-processing steps described in Section 4.1.4, which takes only a few seconds.
@@ A.1.7 Are error slices distinct bugs?
A.2 Generated attributes and tags by HiBug2 and the baseline methods
@@
Table 5: The full list of attributes and tags generated by HiBug2 for the pose estimation task. For tag sets containing more than five tags, the remaining tags are omitted with ellipses
@@
We present a full list of attributes and tags generated by HiBug2 and HiBug for pose estimation in Table 4 and Table 5. The attributes generated by HiBug primarily focus on the main object, often neglecting task-specific requirements and potential errors. Furthermore, the tags frequently lack semantic consistency; for example, the tags associated with the attribute “age” include overlapping terms such as “young”, “teen”, and “12”. Similarly, the tags for “skin tone” combine both race and color categories, resulting in semantic ambiguity. This inconsistency can adversely affect the data slicing process, leading to images with similar semantics being assigned to different slices (e.g., slice “teen” versus slice “young”). In contrast, the attributes and tags generated by HiBug2 are specifically tailored to tasks and errors, ensuring semantic consistency. This makes HiBug2 more effective for model debugging and refinement.